# Boundary conditions matter: On the spectrum of infinite quantum graphs

### v2

@@ 4.1. Symmetrisation techniques for isoperimetric inequalities
We start with isoperimetric inequalities for the lowest positive eigenvalue of both Friedrichs and Neumann realisations of the Laplacian on an infinite metric graph $\mathcal{G}$. Our estimates extend several bounds only known in the case where $\mathcal{G}$ is compact. While the inequalities themselves could also be obtained via an approximation argument using surgery principles generalised to infinite graphs, a suitable compact exhaustion, and Lemma 3.8, we wish to go further and also identify the cases of equality. This requires somewhat more intricate proofs based on symmetrisation and the coarea formula, techniques going back to Friedlander [22] and Band–Lévy [3, 9].
@@
There is equality if and only $\mathcal{G}$ is a star consisting of $k$ (finite or infinite) path subgraphs of length $|\mathcal{G}|/k$ each, glued together at a common vertex.
By an infinite path subgraph we mean a path graph which consists of an infinite number of edges but having finite volume; these are isometrically isomorphic to intervals which are open at one endpoint and closed at the other.
For the case of the Friedrichs realisation we have the equivalent of the Dirichlet version of the theorems of Nicaise and Friedlander, see [51, Théorème 3.2] and [22, Lemma 3], respectively. Here and in the sequel, let $\mathfrak{V}\subset\overline{\mathcal{G}}\subset\mathsf{V}\cap\mathfrak{C}(\mathcal{G})$, and in place of the Friedrichs realisation $\mathcal{H}_{\mathrm{F}}$ consider the more general $\mathcal{H}_{\mathfrak{V}}$ of Remark 2.19imposing Dirichlet conditions on an (essentially arbitrary) subset $\mathfrak{V}$ of the union of the set of all vertices and the set of all ends of $\mathcal{G}$.
@@
Finally, for doubly path connected graphs, we have the following improvements of (4.1) (for the first nontrivial eigenvalue $\mu_{2}$) and (4.2). Their counterparts for finite metric graphs are [3, Theorem 2.1] and [9, Lemma 4.3], respectively.
@@
Recall that a symmetric necklace, see [3, Example 1.7], is a metric graph obtained concatenating (a finite or countably infinite number of) pumpkin graphs, each on two equally long edges. We emphasise that in the infinite case it is allowed to have two ends.
@@
Equality in (4.4) implies that $\mathcal{G}$ is a symmetric necklace such that either (1) $\mathfrak{V}$ consists of one end, equipped with the Friedrichs realisation, or (2) the necklace is finite with a single Dirichlet condition imposed at one of its extremities, or (3) $\mathfrak{V}$ consists of one end, equipped with the Neumann realisation, but there is a single Dirichlet condition at the other extremity of the necklace.
@@
Figure 4.1. The three cases where equality in (4.4) holds. Dirichlet conditions (at vertices or ends) are coloured light grey, standard/Neumann conditions are depicted as filled black circles. (1) In the top graph the Friedrichs realisation leads to a formal Dirichlet condition at the end on the right side of the infinite necklace; (2) the middle graph is a compact necklace with a Dirichlet condition at one end and a standard condition at the other; (3) in the bottom row there is a Dirichlet vertex of degree two at the finite end of the necklace and a Neumann condition at the other end.
@@
The proofs will make use of the following technical result, which is a consequence of the coarea formula (cf. [22, Proof of Lemma 3], [3, Proof of Theorem 2.1] and [9, Proof of Theorem 3.4]). Notationally, given a graph $\mathcal{G}$ and a measurable function $f:\mathcal{G}\to\mathbb{R}$, we denote by
@@
[…]
@@
(and extension by continuity to $x=0,L$). Taking $f\in H^{1}(\mathcal{G})$ to be fixed, we also set $n(t):=\#S_{t}=\#\{x\in\mathcal{G}:f(x)=t\}$, which is a nonnegative integer for almost all $t\in\mathbb{R}$ by Lemma 4.5; and in fact is (again, for almost all $t\in\mathbb{R}$) at least 2 under the assumptions of Theorems 4.3 and 4.4, by an argument essentially based on Menger’s Theorem as in [3, 9]. The following result is now standard; its proof, using the properties established in Lemma 4.5, follows the proof of, e.g., [3, Theorem 2.1] essentially verbatim (see in particular Eq. (3.13) there). We therefore omit it.
@@
As a consequence we obtain the functional inequality of Friedlander [22, Lemma 3] for infinite graphs. We set this up to work simultaneously for both $H^{1}$-functions vanishing at at least one point in $\mathcal{G}$ and functions in $H^{1}_{0}(\mathcal{G})$; the result for the latter will immediately imply Theorem 4.2. Recall that $H_{0}^{1}(\mathcal{G};\{y\})\subset H^{1}(\mathcal{G})$ denotes the subspace of $H^{1}$-functions that vanish at $y$, whether $y$ is a vertex or an end.
@@
In either case, by replacing $f$ by $|f|$ we may assume that $f$ is nonnegative. If $y$ is a vertex, then $f\in H_{0}^{1}(\mathcal{G};\{y\})$; since $f$ is continuous and has a zero in $\mathcal{G}$, $f^{\ast}(0)=0$. It follows from Lemma 4.7, the fact that $n(t)\geq 1$ for all $t\in(m,M)$ by continuity of $f$, and the usual one-dimensional inequality (cf. [22, Eq. (2.7)]) that
@@
If there is equality in (4.10), and hence in every step of (4.11), then $n(t)=1$ for almost all $t\in(m,M)=(0,M)$ and hence for all $t\in(0,M)$; moreover, since up to scalar multiples and reflections $f^{\ast}(x)=\sin(\pi x/2L)$, $f$ can only vanish on a set of measure zero, and hence at most at a single point. That is, $f$ can take on any value between 00 and its maximum on $\overline{\mathcal{G}}$ only once. A standard continuity argument shows that $\mathcal{G}$ must be a path graph, with $f$ vanishing at exactly one of its endpoints. The fact that $f$, considered as a function on $[0,L]$, minimises the Rayleigh quotient of the second derivative on $(0,L)$ with a Dirichlet condition at 00, means that it must be proportional to the first eigenfunction $\sin(\frac{\pi}{2|\mathcal{G}|}\,\cdot\,)$. ∎
@@
Since $\mathcal{G}$ has finite volume, $\mathcal{H}_{\mathfrak{V}}$ has compact resolvent and there exists a positive ground state $\psi_{1}\in H^{1}_{0}(\mathcal{G})\cap C(\overline{\mathcal{G}})\cap\oplus_{\mathsf{e}\in\mathsf{E}}C^{1}([0,\ell_{\mathsf{e}}])$. In particular, the function $\psi_{1}$ can only vanish at $\mathfrak{V}$, where it attains its minimum $m$. Therefore, $\psi_{1}$ is an $H^{1}(\mathcal{G})$-function which takes all values in $[0,M]$ in $\overline{\mathcal{G}}$, where $M:=\max_{x\in\overline{\mathcal{G}}}\psi_{1}(x)$. Lemma 4.7 implies that the Rayleigh quotient of $\psi_{1}$ is at least $\operatorname{ess\,inf}_{t\in[0,M]}n(t)^{2}$ times the Rayleigh quotient of its increasing rearrangement $\psi_{1}^{*}$, which in turn is no smaller than $\left(\frac{\pi}{2|\mathcal{G}|}\right)^{2}$, that is, the lowest eigenvalue of the Laplacian on $[0,|\mathcal{G}|]$ with mixed Dirichlet/Neumann boundary conditions.
@@
We refer to [10, Definition 3.2] for the definition of cutting through a vertex; we note that the definition of cuts originally given for compact metric graphs makes equal sense in the infinite case, since cutting through vertex is a local graph operation and the cutting procedure in Lemma 4.9 is iterative, i.e., we do not need to cut through infinitely many vertices simultaneously. The proof of Lemma 4.9 is analogous to the one of [10, Theorem 3.4] and therefore omitted.
Note that it is always possible to obtain a tree from any locally finite graph $\mathcal{G}$, upon cutting a countable number of times. We will thus assume for the rest of the section that $\mathcal{G}$ is a (locally finite, connected) tree. In what follows, we will need the following notation. For an arbitrary point $x\in\mathcal{G}$ the graph $\mathcal{G}\backslash\{x\}$ consists of $p_{x}\in\mathbb{N}_{0}$ connected components. We denote the closure of these $p_{x}$ components by $\left\{\mathcal{G}^{1}(x),\ldots\mathcal{G}^{p_{x}}(x)\right\}$; these graphs are also trees. Inductively, we can remove a finite number of points $\{x_{1},\ldots x_{n}\}$. The set of the resulting graphs will be denoted by $\mathcal{G}(x_{1},\ldots,x_{n})$, and an element of this set with $k(x_{1},\ldots,x_{n})$ elements by $\mathcal{G}^{j}(x_{1},\ldots x_{n})$ with $1\leq j\leq k(x_{1},\ldots,x_{n})$.
The next lemma is a generalisation of [22, Lemma 4].
@@
[…]
@@
We take $\varepsilon>0$ with $\min(|\mathcal{G}|-l,l)<|\mathcal{G}|-\varepsilon$. We take a compact exhaustion $(\mathcal{G}_{n})_{n\in\mathbb{N}}$ of $\mathcal{G}$ and fix $n$ large enough that $|\mathcal{G}\backslash\mathcal{G}_{n}|<\varepsilon$. Now consider $\mathcal{G}_{n}$. This graph is compact, and $\mathcal{G}\setminus\mathcal{G}_{n}$ consists of at most finitely many (possibly infinite) graphs $\widehat{\mathcal{G}}_{k}$, $k=1,\ldots,m$, attached as pendants to $\mathcal{G}_{n}$ at finitely many vertices (cf. the discussion after Proposition 2.10). The sum of the volumes of these $\widehat{\mathcal{G}}_{k}$ does not exceed $\varepsilon$. We replace each of these $m$ graphs $\widehat{\mathcal{G}}_{k}$ by an edge $\mathsf{e}_{k}$ of length $|\widehat{\mathcal{G}}_{k}|$. The resulting graph, call it $\widehat{\mathcal{G}}_{n}$, satisfies the conditions of [22, Lemma 4]. Thus, we can choose the point $x$ from this lemma. This point $x$ cannot be an element of the interior of an edge $\mathsf{e}_{k}$, since otherwise one of the resulting graphs would have volume exceeding $|\mathcal{G}|-\varepsilon$. Since $x$ cannot be a vertex of degree one, we conclude that $x\in\mathcal{G}_{n}$. Per construction $x$ also has the desired properties in the graph $\mathcal{G}$. ∎
The following generalises [22, Lemma 2]; the proof in the compact case may be repeated verbatim and is thus omitted.
@@
Let $\mathcal{G}$ be a locally finite metric tree with finite volume $|\mathcal{G}|>0$ and let $n\in\mathbb{N}$ with $n\geq 2$. Then there exist $n-1$ points $x_{1},\ldots x_{n-1}\in\mathcal{G}$ such that, for any element of the set $\mathcal{G}(x_{1},\ldots,x_{n-1})$, we have
@@
We can finally give the proof of Theorem 4.1, which closely follows that of [22, Theorem 1].
@@
We still have to show the case of equality. The only difference to the proof for the finite case is that instead of segments with a Dirichlet endpoint $y$ we may also have infinite path graphs of finite length, with a Dirichlet endpoint $y$. Note that at the graph end of any such path the same Neumann condition holds as for a leaf of a finite path graph. Applying in particular Lemma 4.11, the rest of the proof carries over verbatim from [22, Section 2], so we omit it. ∎
@@ 4.2.1. Diameter
Apart from the volume, an important quantity in spectral geometry is the diameter. It was proved in [36, Section 5] that neither lower nor upper bounds on the spectral gap are generally possible in terms of diameter alone. However, estimates can be obtained if the diameter is complemented by volume, or if the graph has a special topological structure. In what follows we extend such estimates to infinite graphs.
We start with a bound in terms of volume and diameter. In the case of compact metric graphs, the following theorem is found in [53, Theorem 4.4.6]; see also [35, Theorem 1.1] and [36, Theorem 7.2] for earlier iterations.
@@
Choose a compact exhaustion of $\mathcal{G}$, and apply [53, Theorem 4.4.6]. Proposition 2.12 and Lemma 3.8(2) then yield (4.15) on $\mathcal{G}$. ∎
One can, similarly, generalise the result of [35, Theorem 1.2], which gives a lower bound on $\mu_{k}$ in terms of $|\mathcal{G}|$ and $\operatorname{Diam}(\mathcal{G})$, to locally finite graphs with finite Betti number. This estimate can be greatly improved if $\mathcal{G}$ is a (finite or infinite) tree. In this case, we obtain:
@@
Note that due to Proposition 3.2, the assumptions of Proposition 4.14 imply that the spectrum of $\mathcal{H}_{\mathfrak{V}}$ is purely discrete. The proof of Proposition 4.14 follows precisely along the lines of [9, Lemma 4.6] and is therefore omitted.
@@
Other lower bounds on $\lambda_{1}(\mathcal{G})$ are obtained in [59, Corollary 3.6 and Theorem 4.1.(i)] in terms of both diameter and further related quantities (the height and reduced height of the tree), but under the additional assumption that the tree is radially symmetric and a Dirichlet condition is imposed on its root.
The discussion following [14, Proposition 2.4] shows that the assumption that $\mathcal{G}$ is a tree cannot generally be dropped.

### v3

@@ 4.1. Symmetrisation techniques for isoperimetric inequalities
We start with isoperimetric inequalities for the lowest positive eigenvalue of both Friedrichs and Neumann realisations of the Laplacian on an infinite metric graph $\mathcal{G}$. Our estimates extend several bounds only known in the case where $\mathcal{G}$ is compact. While the inequalities themselves could also be obtained via an approximation argument using surgery principles generalised to infinite graphs, a suitable compact exhaustion, and Lemma 3.9, we wish to go further and also identify the cases of equality. This requires somewhat more intricate proofs based on symmetrisation and the coarea formula, techniques going back to Friedlander [27] and Band–Lévy [8, 14]. Indeed, the following extends [27, Theorem 1] to the case of graphs with nonempty set of ends; its proof will be presented at the end of this subsection, after collecting a number of auxiliary results.
@@
There is equality if and only $\mathcal{G}$ is a star consisting of $k$ path subgraphs of length $|\mathcal{G}|/k$ each, glued together at a common vertex.
For the case of the Friedrichs realisation we have the equivalent of the Dirichlet version of the theorems of Nicaise and Friedlander, see [55, Théorème 3.2] and [27, Lemma 3], respectively. Here and in the sequel, let $\mathfrak{V}\subset\overline{\mathcal{G}}\subset\mathsf{V}\cap\mathfrak{C}(\mathcal{G})$, and in place of the Friedrichs realisation $\mathcal{H}_{\mathrm{F}}$ consider the more general $\mathcal{H}_{\mathfrak{V}}$ of Remark 2.20imposing Dirichlet conditions on an (essentially arbitrary) subset $\mathfrak{V}$ of the union of the set of all vertices and the set of all ends of $\mathcal{G}$.
@@
Finally, for doubly path connected graphs, we have the following improvements of (4.1) (for the first nontrivial eigenvalue $\mu_{2}$) and (4.2). Their counterparts for finite metric graphs are [8, Theorem 2.1] and [14, Lemma 4.3], respectively.
@@
Recall that a symmetric necklace, see [8, Example 1.7], is a metric graph obtained concatenating (a finite or countably infinite number of) pumpkin graphs, each on two equally long edges. We emphasise that in the infinite case it is allowed to have two ends.
@@
Equality in (4.4) implies that $\mathcal{G}$ is a symmetric necklace such that either (1) $\mathfrak{V}$ consists of one end, equipped with the Friedrichs realisation, or (2) the necklace is finite with a single Dirichlet condition imposed at one of its extremities, or (3) $\mathfrak{V}$ consists of one end, equipped with the Neumann realisation, but there is a single Dirichlet condition at the other extremity of the necklace, or (4) the necklace has two ends, with a Neumann realisation at one and a Dirichlet realisation at the other.
@@
Figure 4.1. The four cases where equality in (4.4) holds. Dirichlet conditions (at vertices or ends) are coloured light grey, standard/Neumann conditions are depicted as filled black circles. (1) In the top graph the Friedrichs realisation leads to a formal Dirichlet condition at the end on the right side of the infinite necklace; (2) the second graph is a compact necklace with a Dirichlet condition at one end and a standard condition at the other; (3) in the third row there is a Dirichlet vertex of degree two at the finite end of the necklace and a Neumann condition at the other end; (4) the bottom graph has a Dirichlet condition at the end on the left and a Neumann condition at the end on the right.
@@
The proofs will make use of the following technical result, which is a consequence of the coarea formula (cf. [27, Proof of Lemma 3], [8, Proof of Theorem 2.1] and [14, Proof of Theorem 3.4]). Notationally, given a graph $\mathcal{G}$ and a measurable function $f:\mathcal{G}\to\mathbb{R}$, we denote by
@@
Let $\mathcal{G}$ be any metric graph on a countable edge set, let $f\in C(\mathcal{G})\cap\bigoplus_{\mathsf{e}\in\mathsf{E}}C^{1}([0,\ell_{\mathsf{e}}])$ and let $\varphi\in L^{1}_{\textrm{loc}}(\mathcal{G})$ be nonnegative. Then, with the notation introduced above,
@@
(and extension by continuity to $x=0,L$). Taking $f\in H^{1}(\mathcal{G})$ to be fixed, we also set $n(t):=\#S_{t}=\#\{x\in\mathcal{G}:f(x)=t\}$, which is a nonnegative integer for almost all $t\in\mathbb{R}$ by Lemma 4.5; and in fact is (again, for almost all $t\in\mathbb{R}$) at least 2 under the assumptions of Theorems 4.3 and 4.4, by an argument essentially based on Menger’s Theorem as in [8, 14]. The following result is now standard; its proof, using the properties established in Lemma 4.5, follows the proof of, e.g., [8, Theorem 2.1] essentially verbatim (see in particular Eq. (3.13) there). We therefore omit it.
@@
As a consequence we obtain the functional inequality of Friedlander [27, Lemma 3] for infinite graphs. We set this up to work simultaneously for both $H^{1}$-functions vanishing at at least one point in $\mathcal{G}$ and functions in $H^{1}_{0}(\mathcal{G})$; the result for the latter will immediately imply Theorem 4.2. Recall that $H_{0}^{1}(\mathcal{G};\{y\})\subset H^{1}(\mathcal{G})$ denotes the subspace of $H^{1}$-functions that vanish at $y$, whether $y$ is a vertex or an end.
@@
In either case, by replacing $f$ by $|f|$ we may assume that $f$ is nonnegative. If $y$ is a vertex, then $f\in H_{0}^{1}(\mathcal{G};\{y\})$; since $f$ is continuous and has a zero in $\mathcal{G}$, $f^{\ast}(0)=0$. It follows from Lemma 4.7, the fact that $n(t)\geq 1$ for all $t\in(m,M)$ by continuity of $f$, and the usual one-dimensional inequality (cf. [27, Eq. (2.7)]) that
@@
If there is equality in (4.10), and hence in every step of (4.11), then $n(t)=1$ for almost all $t\in(m,M)=(0,M)$ and hence for all $t\in(0,M)$; moreover, since up to scalar multiples and reflections $^{\ast}(x)=\sin(\pi x/2L), ff$ can only vanish on a set of measure zero, and hence at most at a single point. That is, $f$ can take on any value between 0 and its maximum on $\overline{\mathcal{G}}$ only once. A standard continuity argument shows that $\mathcal{G}$ must be a path graph, with $f$ vanishing at exactly one of its endpoints. The fact that $f$, considered as a function on $[0,L]$, minimises the Rayleigh quotient of the second derivative on $(0,L)$ with a Dirichlet condition at 0, means that it must be proportional to the first eigenfunction $\sin(\frac{\pi}{2|\mathcal{G}|}\,\cdot\,)$. ∎
@@
Since $\mathcal{G}$ has finite volume, $\mathcal{H}_{\mathfrak{V}}$ has compact resolvent and there exists a positive ground state $\psi_{1}\in H^{1}_{0}(\mathcal{G})\cap C(\overline{\mathcal{G}})\cap\oplus_{\mathsf{e}\in\mathsf{E}}C^{1}([0,\ell_{\mathsf{e}}])$. In particular, the function $\psi_{1}$ can only vanish at $\mathfrak{V}$, where it attains its minimum $m$. Therefore, $\psi_{1}$ is an $H^{1}(\mathcal{G})$-function which takes all values in $[0,M]$ in $\overline{\mathcal{G}}$, where $M:=\max_{x\in\overline{\mathcal{G}}}\psi_{1}(x)$. Lemma 4.7 implies that the Rayleigh quotient of $\psi_{1}$ is at least $\operatorname{ess\,inf}_{t\in[0,M]}n(t)^{2}$ times the Rayleigh quotient of its increasing rearrangement $\psi_{1}^{*}$, which in turn is no smaller than $\left(\frac{\pi}{2|\mathcal{G}|}\right)^{2}$, that is, the lowest eigenvalue of the Laplacian on $0,|\mathcal{G}|]$ with mixed Dirichlet/Neumann boundary conditions.
@@
We refer to [15, Definition 3.2] for the definition of cutting through a vertex; we note that the definition of cuts originally given for compact metric graphs makes equal sense in the infinite case, since cutting through vertex is a local graph operation and the cutting procedure in Lemma 4.9 is iterative, i.e., we do not need to cut through infinitely many vertices simultaneously. The proof of Lemma 4.9 is analogous to the one of [15, Theorem 3.4] and therefore omitted.
[…]
The next lemma is a generalisation of [27, Lemma 4].
@@
$|\mathcal{G}^{1}(x)|\leq|\mathcal{G}|-l\quad\text{and}\quad|\mathcal{G}^{i}(x)|\leq l\qquad\text{for all}\quad 2\leq i\leq p_{x}.$
@@
We take $\varepsilon>0$ with $\min(|\mathcal{G}|-l,l)<|\mathcal{G}|-\varepsilon$. We take a compact exhaustion $(\mathcal{G}_{n})_{n\in\mathbb{N}}$ of $\mathcal{G}$ and fix $n$ large enough that $|\mathcal{G}\backslash\mathcal{G}_{n}|<\varepsilon$. Now consider $\mathcal{G}_{n}$. This graph is compact, and $\mathcal{G}\setminus\mathcal{G}_{n}$ consists of at most finitely many (possibly infinite) graphs $\widehat{\mathcal{G}}_{k}$, $k=1,\ldots,m$, attached as pendants to $\mathcal{G}_{n}$ at finitely many vertices (cf. the discussion after Proposition 2.10). The sum of the volumes of these $\widehat{\mathcal{G}}_{k}$ does not exceed $\varepsilon$. We replace each of these $m$ graphs $\widehat{\mathcal{G}}_{k}$ by an edge $\mathsf{e}_{k}$ of length $|\widehat{\mathcal{G}}_{k}|$. The resulting graph, call it $\widehat{\mathcal{G}}_{n}$, satisfies the conditions of [27, Lemma 4]. Thus, we can choose the point $x$ from this lemma. This point $x$ cannot be an element of the interior of an edge $\mathsf{e}_{k}$, since otherwise one of the resulting graphs would have volume exceeding $|\mathcal{G}|-\varepsilon$. Since $x$ cannot be a vertex of degree one, we conclude that $x\in\mathcal{G}_{n}$. Per construction $x$ also has the desired properties in the graph $\mathcal{G}$. ∎
The following generalises [27, Lemma 2]; the proof in the compact case may be repeated verbatim and is thus omitted.
@@
Let $\mathcal{G}$ be a locally finite metric tree with finite volume $|\mathcal{G}|>0$ and let $n\in\mathbb{N}$ with $n\geq 2$. Then there exist n1n-1 points $x_{1},\ldots x_{n-1}\in\mathcal{G}$ such that, for any element of the set $\mathcal{G}(x_{1},\ldots,x_{n-1})$, we have
@@
We can finally give the proof of Theorem 4.1, which closely follows that of [27, Theorem 1].
@@
We still have to show the case of equality. The only difference to the proof for the finite case is that instead of segments with a Dirichlet endpoint $y$ we may also have infinite path graphs of finite length, with a Dirichlet endpoint $y$. Note that at the graph end of any such path the same Neumann condition holds as for a leaf of a finite path graph. Applying in particular Lemma 4.11, the rest of the proof carries over verbatim from [27, Section 2], so we omit it. ∎
@@ 4.2.1. Diameter
Apart from the volume, an important quantity in spectral geometry is the diameter. It was proved in [41, Section 5] that neither lower nor upper bounds on the spectral gap are generally possible in terms of diameter alone. However, estimates can be obtained if the diameter is complemented by volume, or if the graph has a special topological structure. In what follows we extend such estimates to infinite graphs.
We start with a bound in terms of volume and diameter. In the case of compact metric graphs, the following theorem is found in [58, Theorem 4.4.6]; see also [40, Theorem 1.1] and [41, Theorem 7.2] for earlier iterations.
@@
Choose a compact exhaustion of $\mathcal{G}$, and apply [58, Theorem 4.4.6]. Proposition 2.13 and Lemma 3.9(2) then yield (4.15) on $\mathcal{G}$. ∎
One can, similarly, generalise the result of [40, Theorem 1.2], which gives a lower bound on $\mu_{k}$ in terms of $|\mathcal{G}|$ and $\operatorname{Diam}(\mathcal{G})$, to locally finite graphs with finite Betti number. This estimate can be greatly improved if $\mathcal{G}$ is a (finite or infinite) tree. In this case, we obtain:
@@
Note that due to Proposition 3.2, the assumptions of Proposition 4.14 imply that the spectrum of $\mathcal{H}_{\mathfrak{V}}$ is purely discrete. The proof of Proposition 4.14 follows precisely along the lines of [14, Lemma 4.6] and is therefore omitted.
@@
Other lower bounds on $\lambda_{1}(\mathcal{G})$ are obtained in [62, Corollary 3.6 and Theorem 4.1.(i)] in terms of both diameter and further related quantities (the height and reduced height of the tree), but under the additional assumption that the tree is radially symmetric and a Dirichlet condition is imposed on its root.
The discussion following [19, Proposition 2.4] shows that the assumption that $\mathcal{G}$ is a tree cannot generally be dropped.